Blog from May, 2017


As part of the Laboratory’s ongoing efforts to improve cyber security, TREX will require VPN for off site access beginning June 29th 2017.    Users can get more information about VPN installation at and get assistance at  VPN service is provided at no charge for all LBL employees and affiliates.

Limiting access to the TREX application to those on our network or using VPN helps to limit the exposure of the application to hostile attackers on the internet.  Applications exposed on the internet are subject to constant attack, and while LBL’s security systems help to reduce this exposure, IT judged that further reducing this risk was appropriate for TREX at this time.

If you have questions about VPN or feedback about this change, please visit

Note that TREX is also undergoing a technical upgrade at the same time.  For more details, visit

Creative Services Joins IT

As part of the broader transformation of Public Affairs into Strategic Communications, IT is pleased to welcome the staff of Creative Services to the Division.   We look forward to deeper integration between our teams in the coming months.  All of CSO's services from printing to design to photography, video, and web development remain available to the Lab community via the same contacts:  and

The following staff have joined IT:

Cait Youngquist
Ivan Berry
Phil Butler
Theresa Duque
Ross Lyon
Marilyn Chung
Susan Brand
Zofia Rostoman
Jon Rhodin

The Indico software on has been upgraded from v1.2 to v1.2.2rc1 to fix security vulnerabilities.


May 10 9:30AM All user-facing systems except for scientific unix/scientific clusters are believed to be fully available.

May 10 8:24AM Restoration activities are ongoing.  Some business systems which were restored need to be taken down to address a hardware issue (including FMS).   This outage should take < 1 hr. Telephony and networking appear to be working correctly - please report issues to   There is no ETA for scientific computing/clusters but restoration is in progress and no hardware failures have yet been identified.

The power outage which occurred ~9:30APM May 9th has been partially resolved.  Power has returned to most areas of the site (see Level 1 emails).   Most IT systems are operational, but some are still being restored.    Scientific Computing/Clusters/HPCS have no ETA for restoration at this time.


There is a power outage impacting numerous systems on the site as of 9:40PM May 9 2017.   IT is investigating.   No estimated return to service.